OpenStudy (anonymous):

find all the roots and check your answer. x^2 + 5x - 14 = 0

OpenStudy (anonymous):

Do you know how to do trinomial factoring?

OpenStudy (anonymous):

is that with the formula x^2 + bx + c?

OpenStudy (anonymous):

No. For example, you can factor x^2-9x+20 to be (x-4)(x-5). Can you figure out how to factor x^2+5x-14 in the same way?

OpenStudy (anonymous):

okay I think I understand now because 5*4 = 20 and 5+4=9

OpenStudy (anonymous):

Yes, so you want to find two numbers that multiply to get -14 and add to get 5

OpenStudy (anonymous):

okay thank you so much

OpenStudy (anonymous):

Wait you're not done yet! First write it in factored form.

OpenStudy (anonymous):

o okay I know the numbers will be 7 and 2 (x+2) (x-5)

OpenStudy (anonymous):

But you used 2 and 5 :P 7 and 2 is correct, try again!

OpenStudy (anonymous):

i meant 2,7 that's what i wrote on my problem sorry

OpenStudy (anonymous):

(x-2) (x+7)

OpenStudy (anonymous):

Perfect, very good. So a root of an equation means when it equals 0. So now you want to solve: (x-2)(x+7)=0. If either (x-2) OR (x-7) is 0, then the whole thing will equal 0. So you can now write it as two equations and solve them independently: x-2=0 x+7=0

OpenStudy (anonymous):

so x = 2 and x = 7

OpenStudy (anonymous):

-7

OpenStudy (anonymous):

Bravo! That's it! So that's how you solve for the roots of a quadratic equation :) If you can't factor it, you have to use the "quadratic formula" but I'm not sure you've learned it yet.

OpenStudy (anonymous):

thank you